# Maximum Likelihood Estimation Based Complex-Valued Robust Chinese Remainder Theorem and Its Fast Algorithm

### v1

@@ VI Conclusion
This paper proposes an efficient C-CRT algorithm based on MLE, enabling robust reconstruction of complex numbers from erroneous remainders. The optimal common remainder can be determined using $L$ searches in both real and imaginary parts. Additionally, a necessary and sufficient condition for the C-CRT algorithm to achieve robust reconstruction is provided. In simulations, all the theoretical results are verified and it is illustrated that the proposed algorithm outperforms the current two-stage CRT and has been successfully applied to ADCs for complex-valued bandlimited signals.
@@ VI-A Proof of Properties 1 to 7
Property 1: Since $\mathsf{z}\in\mathcal{F}_{\mathsf{M}}$, we can express $\mathsf{z}$ as $\rho e^{\mathrm{i}\theta}(a+b\mathrm{i})$ by (2), where $10\leq a,b<$ 1. Thus, $\mathsf{z}e^{-\mathrm{i}\theta}=\rho(a+b\mathrm{i})$. According to (2), we have $\mathsf{z}e^{-\mathrm{i}\theta}\in\mathcal{F}_{\rho}$.
Property 2: For any point $\mathsf{z}$, multiplying it by $e^{-\mathrm{i}\theta}$ results in a rotation of $\mathsf{z}$ by $-\theta$. By Property 1, we know that the region $\mathcal{F}_{\mathsf{M}}$ is rotated to $\mathcal{F}_{\mathsf{\rho}}$. Since $\mathcal{F}_{\mathsf{\rho}}$ is a square with side length $\rho$, we have that $\mathcal{F}_{\mathsf{M}}$ is also a square and the area of $\mathcal{F}_{\mathsf{M}}$ is $\rho^{2}$.
[…]
Property 4: Since $[\mathsf{z}+\mathsf{k}]=[\mathsf{z}]+\mathsf{k}$ holds for any complex number $\mathsf{z}$, we obtain by the definition of the circular distance in (16) that $d_{\mathsf{M}}(\mathsf{x}+\mathsf{kM},\mathsf{y})=d_{\mathsf{M}}(\mathsf{x},\mathsf{y})$ and $d_{\mathsf{\mathsf{M}}}(\mathsf{x},\mathsf{y}+\mathsf{kM})=d_{\mathsf{\mathsf{M}}}(\mathsf{x},\mathsf{y})$. Furthermore, it follows from (5) that $d_{\mathsf{M}}(\mathsf{x},\mathsf{y})=d_{\mathsf{M}}\left(\mathsf{x},\mathsf{y}-\mathsf{M}\left\lfloor\frac{\mathsf{y}}{\mathsf{M}}\right\rfloor\right)=d_{\mathsf{M}}(\mathsf{x},\left\langle\mathsf{y}\right\rangle_{\mathsf{M}})$.
Property 5: Let $\mathsf{x}-\mathsf{y}=\mathsf{M}(c+d\mathrm{i})$, where $-\frac{1}{2}\leq c,d<\frac{1}{2}$. Note that $\left[\frac{\mathsf{M}(c+d\mathrm{i})}{\mathsf{M}}\right]=0$. Hence, $d_{\mathsf{M}}(\mathsf{x},\mathsf{y})=\mathsf{x}-\mathsf{y}$.
Property 6: Let $\mathsf{x}-\mathsf{y}=\mathsf{M}(a+b\mathrm{i})$. Then, we have either $\mathsf{x}-\mathsf{y}\in\mathcal{S}_{\mathsf{M}}$ or $\mathsf{x}-\mathsf{y}\not\in\mathcal{S}_{\mathsf{M}}$. For the first case, we have three subcases: 1) $-\frac{1}{2}<a<\frac{1}{2},b=-\frac{1}{2}$; 2) $-\frac{1}{2}<b<\frac{1}{2}$, $a=-\frac{1}{2}$; 3) $a=b=-\frac{1}{2}$. By Property 5, we have $|d_{\mathsf{M}}(\mathsf{x},\mathsf{y})|=|\mathsf{x}-\mathsf{y}|$ for these three subcases. For the second case, we have three subcases: 1) $-\frac{1}{2}\leq a<\frac{1}{2}$ and $b=\frac{1}{2}$; 2) $-\frac{1}{2}\leq b<\frac{1}{2}$ and $a=\frac{1}{2}$; 3) $a=b=\frac{1}{2}$. Without loss of generality, we only prove subcase 1). Since $\left[\frac{\mathsf{M}(a+b\mathrm{i})}{\mathsf{M}}\right]=\mathrm{i}$, we have $|d_{\mathsf{M}}(\mathsf{x},\mathsf{y})|=|\mathsf{M}(a+b\mathrm{i})-\mathsf{M}\mathrm{i}|=|\mathsf{M}(a-\frac{1}{2}\mathrm{i})|=|\mathsf{M}(a+\frac{1}{2}\mathrm{i})|=|\mathsf{x}-\mathsf{y}|$.
@@ VI-B Proof of Theorem 5
[…]
@@ VI-C Proof of Theorem 6
The sufficiency has been proven, now we prove the necessity. If not, there must exists a non-empty set $\mathcal{V}_{1}\subset\{\Delta\mathsf{r}_{1},\Delta\mathsf{r}_{2},\ldots,\Delta\mathsf{r}_{L}\}$ satisfying $\sum_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta\mathsf{r}_{j}\in\mathcal{V}_{1}}w_{j}}-\sum_{\Delta\mathsf{r}_{i}\in\overline{\mathcal{V}}_{1}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta\mathsf{r}_{j}\in\overline{\mathcal{V}}_{1}}w_{j}}<-\frac{M}{2}$. Then we have
$\sum\limits_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}\left(\mathrm{Re}(\Delta\mathsf{r}_{i}-\overline{\Delta\mathsf{r}})\right)<-\frac{M}{2}\Big{(}1-\sum\limits_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}\Big{)}\sum\limits_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}.$
Let $\tilde{\mathsf{r}}^{c}=\left\langle\mathsf{r}^{c}+\overline{\Delta\mathsf{r}}+M\sum_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}\right\rangle_{M}$, then we can obtain
$\sum_{i=1}^{L}w_{i}d_{M}^{2}\big{(}\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M}\big{)}$
$\hskip 56.9055pt>\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Re}(\tilde{\mathsf{r}}^{c})\right).$
This leads to $\mathrm{Re}(\hat{\mathsf{r}}^{c})\neq\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M}$.
On the other hand, we obtain from $\hat{\mathsf{N}}-\mathsf{N}=\overline{\Delta\mathsf{r}}$ that $M\mathsf{N}_{0}+\hat{\mathsf{r}}^{c}-M\mathsf{N}_{0}-\mathsf{r}^{c}=\overline{\Delta\mathsf{r}}$. Hence, $\hat{\mathsf{r}}^{c}=\langle\mathsf{r}^{c}+\overline{\Delta\mathsf{r}}\rangle_{M}$. Consequently, $\mathrm{Re}(\hat{\mathsf{r}}^{c})=\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M}$, which leads to a contradiction. ∎

### v3

@@ VI Conclusion
This paper proposes an efficient C-CRT algorithm based on MLE, enabling robust reconstruction of complex numbers from erroneous remainders modulo several Gaussian integers. The optimal common remainder can be determined using $L$ searches in both real and imaginary parts. Additionally, a necessary and sufficient condition for the C-CRT algorithm to achieve robust reconstruction is provided. In simulations, all the theoretical results are verified and it is illustrated that the proposed algorithm outperforms the current two-stage CRT and has been successfully applied to multi-channel SR-ADCs for complex-valued bandlimited signals.
@@ VI-A Proof of Properties 1 to 7
Property 1: Since $\mathsf{z}\in\mathcal{F}_{\mathsf{M}}$, we can express $\mathsf{z}$ as $\rho e^{\mathrm{i}\theta}(a+b\mathrm{i})$ by (2), where $0\leq a,b<1$. Thus, $\mathsf{z}e^{-\mathrm{i}\theta}=\rho(a+b\mathrm{i})$. According to (2), we have $\mathsf{z}e^{-\mathrm{i}\theta}\in\mathcal{F}_{\rho}$.
Property 2: For any point $\mathsf{z}$, multiplying it by $e^{-\mathrm{i}\theta}$ results in a rotation of $\mathsf{z}$ by $\theta$. By Property 1, we know that the region $\mathcal{F}_{\mathsf{M}}$ is rotated to $\mathcal{F}_{\mathsf{\rho}}$. Since $\mathcal{F}_{\mathsf{\rho}}$ is a square with side length $\rho$, we have that $\mathcal{F}_{\mathsf{M}}$ is also a square and the area of $\mathcal{F}_{\mathsf{M}}$ is $\rho^{2}$.
Property 3: Let $\mathsf{x}-\mathsf{y}=\mathsf{M}(a+b\mathrm{i})$. By (17), we have $\begin{aligned} d_{\mathsf{M}}(\mathsf{x},\mathsf{y})=\mathsf{M}(a-[a])+\mathsf{M}(b-\left[b\right])\mathrm{i}.\end{aligned}$ Since $\frac{1}{2}\leq a-[a]<\frac{1}{2}$ and $\frac{1}{2}\leq b-[b]<\frac{1}{2}$, we have $d_{\mathsf{M}}(\mathsf{x},\mathsf{y})\in\mathcal{S}_{\mathsf{M}}$.
Property 4: Since $\mathsf{z}+\mathsf{k}]=[\mathsf{z}]+\mathsf{k}$ holds for any complex number $\mathsf{z}$, we obtain by the definition of the circular distance in (17) that $d_{\mathsf{M}}(\mathsf{x}+\mathsf{kM},\mathsf{y})=d_{\mathsf{M}}(\mathsf{x},\mathsf{y})$ and $d_{\mathsf{\mathsf{M}}}(\mathsf{x},\mathsf{y}+\mathsf{kM})=d_{\mathsf{\mathsf{M}}}(\mathsf{x},\mathsf{y})$. Furthermore, it follows from (5) that
$d_{\mathsf{M}}(\mathsf{x},\mathsf{y})=d_{\mathsf{M}}\left(\mathsf{x},\mathsf{y}-\mathsf{M}\left\lfloor\frac{\mathsf{y}}{\mathsf{M}}\right\rfloor\right)=d_{\mathsf{M}}(\mathsf{x},\left\langle\mathsf{y}\right\rangle_{\mathsf{M}}).$
Property 5: Let $\mathsf{x}-\mathsf{y}=\mathsf{M}(c+d\mathrm{i})$, where $\frac{1}{2}\leq c,d<\frac{1}{2}$. Note that $\left[\frac{\mathsf{M}(c+d\mathrm{i})}{\mathsf{M}}\right]=0$. Hence, $d_{\mathsf{M}}(\mathsf{x},\mathsf{y})=\mathsf{x}-\mathsf{y}$.
Property 6: Let $\mathsf{x}-\mathsf{y}=\mathsf{M}(a+b\mathrm{i})$. Then, we have either $\mathsf{x}-\mathsf{y}\in\mathcal{S}_{\mathsf{M}}$ or $\mathsf{x}-\mathsf{y}\not\in\mathcal{S}_{\mathsf{M}}$. For the first case, we have three subcases: 1) $\frac{1}{2}<a<\frac{1}{2},b=-\frac{1}{2}$; 2) $\frac{1}{2}<b<\frac{1}{2}$, $a=-\frac{1}{2}$; 3) $a=b=-\frac{1}{2}$. By Property 5, we have $|d_{\mathsf{M}}(\mathsf{x},\mathsf{y})|=|\mathsf{x}-\mathsf{y}|$ for these three subcases. For the second case, we have three subcases: 1) $\frac{1}{2}\leq a<\frac{1}{2}$ and $b=\frac{1}{2}$; 2) $\frac{1}{2}\leq b<\frac{1}{2}$ and $a=\frac{1}{2}$; 3) $a=b=\frac{1}{2}$. Without loss of generality, we only prove subcase 1). Since $\left[\frac{\mathsf{M}(a+b\mathrm{i})}{\mathsf{M}}\right]=\mathrm{i}$, we have $|d_{\mathsf{M}}(\mathsf{x},\mathsf{y})|=|\mathsf{M}(a+b\mathrm{i})-\mathsf{M}\mathrm{i}|=|\mathsf{M}(a-\frac{1}{2}\mathrm{i})|=|\mathsf{M}(a+\frac{1}{2}\mathrm{i})|=|\mathsf{x}-\mathsf{y}|$.
@@ VI-B Proof of Theorem 5
By (39), we have
$\frac{M}{2}\leq\sum_{\Delta\mathsf{r}_{i}\in\mathcal{V}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta\mathsf{r}_{j}\in\mathcal{V}}w_{j}}-\sum_{\Delta\mathsf{r}_{i}\in\overline{\mathcal{V}}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta\mathsf{r}_{j}\in\overline{\mathcal{V}}}w_{j}}<\frac{M}{2}.$
According to Theorem 3 in [10], we have
$\mathrm{Re}(\hat{\mathsf{r}}^{c})=\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M}.$
Hence, (41) holds. Similarly, we have
$\mathrm{Im}(\hat{\mathsf{r}}^{c})=\langle\mathrm{Im}(\mathsf{r}^{c})+\mathrm{Im}(\overline{\Delta\mathsf{r}})\rangle_{M}.$
This leads to (42). Since $M\in\mathbb{Z}$, we have
$\hat{\mathsf{r}}^{c}=\langle\mathrm{Re}(\mathsf{r}^{c})+(\mathrm{Re}\overline{\Delta\mathsf{r}})\rangle_{M}+\mathrm{i}\langle\mathrm{Im}(\mathsf{r}^{c})+\mathrm{Im}(\overline{\Delta\mathsf{r}})\rangle_{M}=\langle\mathsf{r}^{c}+\overline{\Delta\mathsf{r}}\rangle_{M}.$
@@ VI-C Proof of Theorem 6
The sufficiency has been proven, now we prove the necessity. Since $\hat{\mathsf{N}}-\mathsf{N}=\overline{\Delta\mathsf{r}}$, we have
$M\hat{\mathsf{N}}_{0}+\hat{\mathsf{r}}^{c}-M\mathsf{N}_{0}-\mathsf{r}^{c}=\overline{\Delta\mathsf{r}}.$
This leads to
$\hat{\mathsf{r}}^{c}=\langle\mathsf{r}^{c}+\overline{\Delta\mathsf{r}}\rangle_{M}.$
Hence,
$\mathrm{Re}(\hat{\mathsf{r}}^{c})=\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M},\ \mathrm{Im}(\hat{\mathsf{r}}^{c})=\langle\mathrm{Im}(\mathsf{r}^{c})+\mathrm{Im}(\overline{\Delta\mathsf{r}})\rangle_{M}.$
If there exists a non-empty set $\mathcal{V}_{1}\subset\mathcal{U}$ satisfying
$\sum_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta\mathsf{r}_{j}\in\mathcal{V}_{1}}w_{j}}-\sum_{\Delta\mathsf{r}_{i}\in\overline{\mathcal{V}}_{1}}\frac{w_{i}\mathrm{Re}(\Delta\mathsf{r}_{i})}{\sum_{\Delta\mathsf{r}_{j}\in\overline{\mathcal{V}}_{1}}w_{j}}\geq\frac{M}{2},$
then we can obtain
$\sum\limits_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}\left(\mathrm{Re}(\Delta\mathsf{r}_{i}-\overline{\Delta\mathsf{r}})\right)\geq\frac{M}{2}\left(1-\sum\limits_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}\right)\sum\limits_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}.$
Let
$\tilde{\mathsf{r}}^{c}=\left\langle\mathsf{r}^{c}+\overline{\Delta\mathsf{r}}-M\sum_{\Delta\mathsf{r}_{i}\in\mathcal{V}_{1}}w_{i}\right\rangle_{M}.$ (48)
Then we have
$\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M}\right)\geq\sum_{i=1}^{L}w_{i}d_{M}^{2}\left(\mathrm{Re}(\tilde{\mathsf{r}}_{i}^{c}),\mathrm{Re}(\tilde{\mathsf{r}}^{c})\right).$ (49)
If the equality of (49) holds, then $\mathrm{Re}(\tilde{\mathsf{r}}^{c})$ is optimal. Hence,
$\mathrm{Re}(\tilde{\mathsf{r}}^{c})=\mathrm{Re}(\hat{\mathsf{r}}^{c})=\langle\mathrm{Re}(\mathsf{r}^{c})+\mathrm{Re}(\overline{\Delta\mathsf{r}})\rangle_{M}.$
By (48), we can obtain that either $\mathcal{V}_{1}=\emptyset$ or $\mathcal{V}_{1}=\mathcal{U}$ holds, which is a contradiction. If the inequality of (49) holds, then $\mathrm{Re}(\hat{\mathsf{r}}^{c})$ is not optimal, which is a contradiction. ∎